robertbays56:

betty gets paid 25 cent per paper to deliver newspapers in her neighborhood when betty is sick she pays her brother bobby 10 cent per paper to deliver them for and if bobby makes 7.80 filling in for betty, how much money does betty get to keep for that day

Madiron:

ok so the first step is to set up a problem

Madiron:

First part of the problem:25-10=? Second part of the problem:7.80*0.15=?

IhelpUuHELPme:

I'm not sure where you got 0.15 from considering the fact tht I've been trying to figure out where that came from- AnYwAyS What we need to find out is the number of papers bobby delivered for betty In words, the total amount that bobby made divided by 10 cents/paper would equal thwe number of papers he delivered \[\frac{ 780 }{ 10 } = 78\] Papers Delivered Now, you have to multiply 25 by 78 | \[25 \times 78 = 1950\] Cents which is Betty's pay for delivering 78 papers. Now, since betty pade bobby $7.80, you would have to find the difference \[1950 - 780 = 1170\] Your final answer would be...Betty gets to keep $11.70 for the day. SImplifie Explanation: What we did was we did the whole equation in cents and then converted them to dollars. \[780 \div 10 = 78\] \[25 \times 78 = 1950\] Cents which is Betty's pay for delivering 78 papers. \[1950 - 780 = 1170\] Betty's Pay - what she paid bobby 1170 converted to dollars is $11.70, the amount betty gets to keep for the day.

Madiron:

see what i did equaled the same thing,11.70

Madiron:

wait

Madiron:

idk anymore

IhelpUuHELPme:

Dude, no where in the equation would there be a need to multiply by 0.15 even if it might give you the same answer- Go by the problem- not something else- EVEN THO it gives you the same answer